### v1

# Qualitative Failures of Image Generation Models and Their Application in Detecting Deepfakes

Abstract
The ability of image and video generation models to create photorealistic images has reached unprecedented heights, making it difficult to distinguish between real and fake images in many cases. However, despite this progress, a gap remains between the quality of generated images and those found in the real world. To address this, we have reviewed a vast body of literature from both academic publications and social media to identify qualitative shortcomings in image generation models, which we have classified into five categories. By understanding these failures, we can identify areas where these models need improvement, as well as develop strategies for detecting deep fakes. The prevalence of deep fakes in today’s society is a serious concern, and our findings can help mitigate their negative impact. In order to support research in this field, a collection of instances where models have failed is made available at here.
Keywords— Generative Models, Image and Video Generation, Qualitative Failures, Deepfakes, Image Forensics, Object and Scene Recognition, Neural Networks, Deep Learning
1 Introduction
Generated images, also known as synthetic images, are created by machine learning algorithms or other software programs, while real images are captured by cameras or other imaging devices. Generated images are not real-world representations of a scene or object, but rather computer-generated approximations. As such, they lack the authenticity and realism of real images.
Despite the abundance of anecdotal evidence shared on social media regarding the weaknesses of image generation models, there has yet to be a comprehensive and systematic analysis of these failures. Often, the examples shared by people are selectively chosen to showcase instances in which the models perform well, which may lead to a biased perception of their capabilities, and an overestimation of their effectiveness. While there have been quantitative studies aimed at evaluating and comparing generative models [4, 6], such as the use of metrics like FID [14], these measures can be difficult to interpret and are usually calculated over large datasets, making them unsuitable for determining the authenticity of individual images. Quantitative measures for detecting deep fakes do exist [23], but they are not as easily accessible to the general public as qualitative measures, which are simpler to carry out.
Figure 1: The Fishmarket, Dieppe, 1902 - Camille Pissarro. When observed more closely, it becomes apparent that the faces in the image lack clarity and numerous details are either incorrect or absent, similar to fake images. Although such images may appear authentic at first glance, scrutinizing them thoroughly is crucial to avoid overlooking errors. It is advisable to conduct a detailed examination of each object within the image by zooming in and analyzing its shape, features, location, and interaction with other objects. This approach allows for a more accurate assessment of the image’s authenticity and being free from errors.
As the quality of generated images continues to improve, it is crucial to conduct more in-depth and precise analyses. Thus far, people have been amazed by the ability of synthesized images to approximate natural scenes. When Photoshop was introduced, significant efforts were made to identify manipulated images, and a similar approach is needed for generated images today. It would be beneficial to compile a set of indicators and other resources to aid in detecting deep fakes.
We present a collection of indicators that can be examined in a single image to determine whether it is genuine or manipulated. Overall, we offer five classes of these indicators including Human and Animal Body Parts, Geometry, Physics, Semantics and Logic, as well as Text, Noise, and Details, for both portraits and natural landscapes. The advantage of utilizing qualitative cues is that they are easily accessible and can be utilized by anyone, potentially serving as the initial step in detecting deep fakes.
Generated images can appear realistic when viewed from a distance or at high resolutions, making it difficult to discern them from actual photographs. However, at lower resolutions, nearly all generated images lack distinguishable characteristics that set them apart from real photographs. To illustrate, refer to Figure 1, which depicts a painting by Camille Pissarro featuring intricate details. While the overall image may seem satisfactory, closer inspection reveals several missing details such as distorted facial features.
This study has a dual purpose. Firstly, it aims to explore the differences between generated images and real-world images. Therefore, this research complements studies that propose quantitative approaches for evaluating generative models. Secondly, it aims to examine qualitative methods that can be employed to identify deepfakes and train individuals to become proficient in this task, with the added benefit of systematically organizing this knowledge.
2 Related work
2.1 Quantitative Approaches to Evaluate Generative Models
Quantitative approaches have emerged as a vital tool to evaluate the performance of generative models. These methods rely on quantitative measures to assess how well a model is able to generate realistic data. One commonly used metric is the Inception Score [26], which evaluates the diversity and quality of generated images based on the classification accuracy of a pre-trained classifier. Another popular approach is the Fréchet Inception Distance [14], which uses feature statistics to compare the distribution of generated data with that of real data. Moreover, other metrics such as precision and recall [25] can be used to evaluate the quality of generated samples in specific domains such as vision, text and audio. These quantitative approaches provide a rigorous and objective way to measure the effectiveness of generative models, helping researchers to improve their models and develop more advanced generative techniques. Large-scale datasets have also been used in studies that focus on the qualitative evaluation of generated images (e.g.  [2]). To obtain a thorough analysis of quantitative metrics for evaluating generative models, please refer to the following references [4, 6, 28, 30].
2.2 Deepfake Detection Methods
Deepfake detection has become an essential area of research due to the increasing sophistication of deep learning algorithms that can generate highly realistic fake images, videos, and audio. As a result, numerous deepfake detection methods have been proposed in recent years, ranging from traditional image and video forensic techniques to advanced deep learning-based approaches. These methods can be broadly categorized into two groups: static and dynamic analysis.
Static analysis methods use handcrafted features to distinguish between real and fake images. Examples of static analysis methods include reverse image search, which compares the content of an image to a large database of known images (e.g.  [8]), and error level analysis, which detects inconsistencies in the compression levels of an image [16]. Another method is the use of noise patterns and artifacts, which are common in images and videos captured by digital cameras and can be used to identify forgeries. For instance, the sensor pattern noise in images captured by digital cameras can be used to authenticate images and detect tampering attempts [19]. In addition, traditional forensic techniques such as shadow analysis, lighting analysis, and perspective analysis can also be used to identify inconsistencies in the shadows, lighting, and perspectives of images.
On the other hand, dynamic analysis methods rely on deep neural networks to analyze the temporal features of video and audio data to detect deepfakes. These methods aim to exploit the fact that deepfakes lack the natural temporal variations and correlations that are present in real videos and audios. For instance, the use of convolutional neural networks (CNNs) has been proposed to detect deepfakes by analyzing the spatial features of images and videos (e.g.  [1, 21, 22, 10]). Similarly, recurrent neural networks (RNNs) have been proposed to analyze the temporal features of video and audio data to detect deepfakes [13]. Moreover, Generative Adversarial Networks (GANs) [12] have been used to generate fake images and videos, but can also be used to detect them by identifying inconsistencies in the generator’s output [18].
Overall, deepfake detection is a challenging problem due to the rapid evolution of deep learning algorithms that can generate more realistic fake content [9]. Thus, a combination of static and dynamic analysis approaches is necessary to achieve effective detection of deepfakes. Additionally, extensive evaluation and comparison of deepfake detection methods are essential to identify their effectiveness and limitations and to guide future research in this area. To read more about this subject, you may want to consult [11, 24, 23, 27] which offer comprehensive reviews on the topic.
3 Qualitative Failures of Image Generation Models
We compiled a list of qualitative failures by examining images from various sources including social media websites such as Twitter, LinkedIn, Discord, and Reddit [^1] A few of the images used in this work were obtained with the consent of a Reddit user named Kronzky., as well as images from the DiffusionDB dataset [29]. These images have been generated by notable generative models such as DALL-E 2, Midjourney, StableDiffusion, and Bing Image Creator. Additionally, we analyzed images from websites such as thisxdoesnotexist.com, whichfaceisreal.com, the Adobe Stock library, and openart.ai. We made sure that the text prompts used to generate images were not intentionally seeking peculiar images. Finally, we manually reviewed the images and filtered out the ones without problems.
3.1 Human and Animal Body Parts
Faces. Since the initial triumphs of GANs, the generation of fake faces has been the most extensively scrutinized category for deep generative models [5]. Faces are comparatively simpler to generate than complex scenes because they are easier to calibrate. In the past, the first generated faces were effortlessly recognizable by humans. However, with the advancement of technology such as StyleGAN [15], the latest examples of generated faces are more challenging to distinguish. Figure 2 illustrates a few faces that were generated with issues. You can evaluate your ability to distinguish between real and computer-generated faces by taking a quiz at whichfaceisreal.com.
Figure 2: Examples of poorly generated faces.
Image Background. When creating deepfakes, issues with the background of the images may arise, particularly in cases where the face is in focus while the surrounding clues are incorrect. The neural network used to generate the images focuses mainly on the face and may not pay as much attention to the surrounding details. This can lead to strange companions or chaotic forms in the background. Additionally, the objects or people next to the primary person in the image may appear unnatural or “mutant”. Figure 3 displays several instances of failures as examples.
Figure 3: Fake images can be exposed through background cues.
Eyes and Gaze. Deep generative models have largely overcome issues with early fake images such as cross-eyed, uncentered or different sized pupils, different colored irises, and non-round pupils, as shown in examples in Figure 4. Early GANs used to produce pupils that were not circular or elliptical like those found in real human eyes, which can be a clue that an image is fake. Reflections in the eyes can also be used to identify fake images. Other clues include irregularities in pupil shape, although this is not always indicative of a fake image since some diseases can cause such irregularities. See the example shown in the bottom-right panel in Figure 4.
Unnatural gaze direction or unrealistic eye movements may be observed in deepfakes, which can indicate that a machine learning algorithm generated or manipulated the image. Please see Figure 5.
Figure 4: Here are some instances of eyes that were generated poorly. The eye in the bottom right corner is an actual photograph of a patient who has an irregularly shaped pupil. You can refer to this link for more details. This case represents a unique manifestation of a condition known as “cat’s eye Adie-like pupil,” which is considered a warning sign for ICE syndrome.
Figure 5: Here are some examples of images where the gaze direction is problematic. In these images, one eye appears to be looking in a different direction compared to the other, similar to a medical condition called Strabismus in the real world. You can check out https://en.wikipedia.org/wiki/Strabismus for additional information on this topic.
Eyeglasses. Algorithms can struggle to create realistic eyeglasses, with frame structures often differing between the left and right sides, or with one side having an ornament and the other not. Sometimes the frame can appear crooked or jagged. The glasses may partially disappear or blend with the head, and they can be asymmetrical. The view through the lens may also be heavily distorted or illogical, and nose pads may be missing or distorted. Please see Figure 6 for some examples.
Figure 6: Some samples of generated eyeglasses with poor quality.
Teeth. Rendering teeth is a difficult task for AI, which often results in odd or asymmetric teeth. When someone’s teeth appear unusual or crooked, there’s a good chance that the image was generated by AI. Semi-regular repeating details like teeth are difficult for models to generate, causing misaligned or distorted teeth. This problem has also been observed in other domains, such as texture synthesis with bricks. Occasionally, an image may display an excessive number of teeth or teeth with abnormal shapes and colors, and in some instances, there may be an insufficient number of incisors. Please see Figure 7 for some examples.
Figure 7: Examples of poorly generated teeth.
Ear and Earrings. Ears in AI-generated images may exhibit discrepancies such as differences in size, one ear appearing higher or bigger than the other, or missing or partially missing earrings. Additionally, earrings may be randomly shaped or not match visually. If earrings are asymmetrical or have different features such as one having an attached earlobe while the other doesn’t or one being longer than the other, it’s likely that the image has been generated by AI. Examples of poorly generated ears and earrings are shown in Figure 8.
Figure 8: Clues that can reveal fake ears, here through earrings.
Hair and Whiskers. The style of hair can differ greatly, which also means there is a lot of intricate detail to capture. This makes it one of the most challenging aspects for a model to render accurately. The generated images may contain stray strands of hair in unusual places, or the hair may appear too straight or streaked. Occasionally, the image may resemble acrylic smudges from a palette knife or brush. Another issue may be a strange glow or halo around the hair. In some cases, the model may bunch hair in clumps or create random wisps around the shoulders, while also including thick stray hairs on the forehead. Please see Figure 9.
Figure 9: Examples of poorly generated hair.
Skin. Deepfakes can be deficient in delicate details and subtleties found in genuine images, like skin texture, pores, or fine lines on someone’s face. The skin tone in deepfakes may appear unnatural or inconsistent, such as a person’s face appearing too pale or too red. Additionally, deepfakes may lack the presence of noise or grain which exists in real images, giving a sense of texture and realism. Without the presence of noise or grain, deepfake images may seem excessively clean or artificial. Some example failures are shown in Figure 10.
Figure 10: Examples of poorly generated skin, absolutely perfect skin with no pores.
Limbs, Hands, and Fingers. The models used for generating deepfakes often fall short when it comes to accurately depicting the intricate details of human extremities. For instance, hands may randomly duplicate, fingers can merge together or there may be too many or too few of them, and third legs may unexpectedly appear while existing limbs may disappear without a trace. Furthermore, limbs may be positioned in unrealistic or impossible poses, or there may be an excess number of them. As a result, deepfakes may exhibit unnatural body language, such as unrealistic gestures or postures that are out of place. See Figs 11 and 12.
Figure 11: Examples of images with poorly generated limbs and distorted body.
Figure 12: Issues with AI-generated fingers.
Clothing. Generative models may produce distorted clothing with various issues, such as asymmetrical, peculiar, or illogical textures or components such as zippers or collars merging with the skin, and textures abruptly changing or ending. Please refe to Figure 13 for some of such failures.
Figure 13: Generating realistic clothing is a challenge for generative models.
3.2 Geometry
Deepfakes may exhibit anomalous or atypical image geometry, with objects appearing to be of an unusual shape or size, in comparison to their expected proportions.
Straight Lines and Edges. AI-generated images may lack the straight lines, seams, and connections found in real-world objects, resulting in wavy, misaligned, and jumpy renderings (e.g. in tiles). Deepfakes can also exhibit inconsistent or unnatural image edges, which refer to the boundaries between different parts of the image. Further, surfaces, which are typically straight, may look somewhat uneven in generated images. Some samples failures are shown in Figure 14.
Figure 14: Examples of lines, edges, and surfaces that are generated poorly by AI.
Perspective. Models lack the ability to understand the 3D world, which results in physically impossible situations when objects cross different planes in a scene. These errors are difficult to detect as our brain often auto-corrects them, requiring a conscious investigation of each angle of the object to identify inconsistencies. Deepfakes can display an unnatural or distorted perspective, where a person’s body appears stretched or compressed unrealistically. Deepfakes may also have inconsistent or unrealistic camera angles, where a person’s face appears to be viewed from an impossible angle or perspective. Some example failures are shown in Figure 15.
Figure 15: Examples of generated images that exhibit issues with perspective.
Symmetry. Due to difficulty managing long-distance dependencies in images, symmetry (reflection, radial, translation, etc) can be challenging for models. For instance, in generated images, eyes may appear heterochromatic and crosseyed, unlike in real life where they tend to point in the same direction and have the same color. Additionally, asymmetry may appear in facial hair, eyeglasses, and the types of collar or fabric used on the left and right sides of clothing. Models may face challenges in maintaining symmetry not only in faces but also in other objects and scenes. For instance, two shoes in a pair or wings in an airplane might not be exactly the same. This is a type of reasoning glitch where the model cannot understand that certain elements should be symmetrical. Some example failures are shown in Figures 16 and 17.
Figure 16: Examples of generated images that display inconsistent symmetry.
Figure 17: Additional examples of generated images that exhibit inconsistent symmetry.
Relative Size. Relative size is a visual perceptual cue that helps us understand the size of objects in relation to one another. It is a powerful cue because it allows us to estimate the size of objects even when we do not have any absolute size reference in the scene. Models, however, fall short in synthesizing objects with objects with sizes proportional to their size in the real world. Some example failures are shown in Figure 18.
Figure 18: Examples of images where there is a violation of relative size.
Other Geometry. Deepfakes exhibit various geometrical anomalies that may reveal their artificiality. For instance, their depth cues can be inconsistent or unnatural, causing the foreground or background to seem blurry or devoid of detail. Moreover, deepfakes often lack parallax, which is the apparent displacement of objects when viewed from different perspectives, resulting in a flat or two-dimensional appearance. Additionally, incorrect or inconsistent motion blur may suggest that certain parts of the image have been manipulated. The absence of occlusion, i.e., the overlapping of objects in the scene, is another telltale sign of deepfakes, as it can make the image look flat or unrealistic. Lastly, deepfakes may display improper image alignment, with objects seeming misaligned or out of place.
3.3 Physics
Generated images that violate physics rules exhibit various cues that can give them away as unrealistic or physically impossible. These cues include objects appearing to float in mid-air without support, shadows that are inconsistent with the light source, reflections or refractions that break the laws of optics, objects passing through each other without interaction, and incorrect physics-based simulations such as fluids or cloth that behave in impossible ways. By identifying these cues, it is possible to identify and distinguish realistic images from those that violate the rules of physics.
Reflection. An effective technique for detecting generated images is to examine the lighting and how it interacts with the elements within the image, and how it causes reflections and shadows. Deepfakes can exhibit artificial reflections that are inconsistent with the natural lighting and environment, such as those in glasses, mirrors, or pupils. The root cause of this issue is that deep generative models lack a proper understanding of reflections. While these models may recognize that an image contains a reflection and typically involves two people (one facing the camera and the other with their back turned), they do not comprehend that the two individuals are, in fact, the same person. Deepfakes may display other lighting effects that do not match real-world environments, such as lens flares, lens distortion, chromatic aberration and unnatural specular highlights. These effects are frequently observed in genuine photographs due to the physical properties of camera lenses and the way light is refracted through them. Some example failures are shown in Figure 19.
Figure 19: Generated images with inconsistent reflections.
Shadow. Deepfakes might not include shadows, which are typically found in real images and contribute to the impression of depth and authenticity. It is important to observe objects without shadows and those with highlights that appear to originate from a different direction than the rest of the image. Additionally, if the photo was taken outdoors in natural light during the afternoon, the setting sun will produce longer shadows than it would at midday, which can be easily identified by scrutinizing the shadow’s length. However, this method may not be as precise in artificial lighting conditions. Finally, if there are multiple objects or people within the scene, their shadows should be consistent with each other. Some generated images with inconsistent shadows are shown in 20.
Figure 20: Generated images with inconsistent shadows.
Objects without Support. When an object or material appears to be floating in mid-air without any visible means of support, it gives the impression that the object is defying gravity or the laws of physics. In reality, all objects are subject to the force of gravity unless they are held up by some other force. When an object appears to be floating, it could be a result of an incorrect rendering or an error in the physics simulation that fails to account for the gravitational force. This type of inconsistency can cause a generated image to look unrealistic or implausible. Some example failures are shown Figure 21.
Figure 21: Generated images where some objects lack visible physical support. Some objects are suspended in mid-air without any explanation or justification. This lack of physical support could result from a failure to properly simulate or model the forces acting on the objects in the scene.
3.4 Semantics and Logic
Images produced by generative models may lack the semantic meaning or contextual relationships present in authentic images. These models tend to focus on the nouns in a given prompt and construct a plausible scene based on them, potentially failing to capture the true relationships between objects. It is crucial to bear in mind that AI lacks an inherent understanding of the world and can only process information in terms of shapes and colors. Complex concepts, such as logical connections and three-dimensional space, are beyond its grasp, resulting in potential difficulties in these areas. For example, when tasked with generating an image of the solar system drawn to scale, a generative model may struggle to maintain the correct planetary order, as demonstrated here.
Spatial Reasoning. Natural scenes are complex and contain a wide range of spatial relationships among objects, such as occlusions, relative distances, and orientations. Capturing these relationships requires the model to have a nuanced understanding of the scene and the objects within it, which can be difficult to achieve without more explicit guidance. Furthermore, some image generation models rely solely on pixel-level reconstruction, without explicitly modeling the underlying semantics or spatial relationships. In these cases, the model may generate images that are visually realistic but lack coherent semantic meaning or accurate spatial relationships among objects. Please see Figure 22 for some examples.
Figure 22: Samples of spatial reasoning from [20]. Images are generated by DALL-E 2 for the following text prompts for columns from left to right: “a red basketball with flowers on it, in front of blue one with a similar pattern”, “a red ball on top of a blue pyramid with the pyramid behind a car that is above a toaster”, “a pear cut into seven pieces arranged in a ring, “In late afternoon in January in New England, a man stands in the shadow of a maple tree”, and “An old man is talking to his parents”.
Context and Scene Composition. Deepfakes can be detected through various inconsistencies such as the background or surroundings not matching the real-world environment, missing contextual details, unnatural object placement, and inconsistent image composition. These irregularities may include inconsistencies in order of objects, missing objects or features, objects appearing in the wrong location or orientation, or unnatural arrangement and placement of objects in the image. Please see Figure 23.
Figure 23: Generated images with problems with context and scene composition.
Other Semantics. Figure 24 depicts several additional generated images that exhibit semantic issues. For instance, one image features a person with his head and feet pointing in opposite directions, while another displays a fragmented pizza that does not cohere into a single entity. In yet another image, a blank painting hangs on the wall, creating a confusing and nonsensical composition.
Figure 24: Additional generated images that exhibit semantic issues.
3.5 Text, Noise, and Details
Text. Generating text and logos in images requires the generative model to understand the relationships between the text and the visual content of the image. This can be challenging because the text and image data have different structures and are not directly aligned with each other. Additionally, text can appear in various locations and orientations within an image, and the context of the text may change depending on the surrounding visual content. Furthermore, generating text that accurately describes the visual content of an image requires a deep understanding of the semantics and context of both the text and the image. While some progress has been made in recent years with the development of methods such as image captioning, it is still an active area of research to develop generative models that can effectively generate text in images. Figure 25 displays instances where the text is incomprehensible. In such cases, the letters appear scrambled or duplicated, and the words are spelled incorrectly.
Figure 25: Generative images that exhibit issues or inconsistencies with the text.
Noise, Color, and Blur Artifacts. Digital distortion in the form of pixelation or imperfect coloring can be present in generated images, particularly around the image edges. Monochrome areas may display semi-regular noise with horizontal or vertical banding, potentially due to the network attempting to replicate cloth textures. Older GANs tend to produce a more noticeable checkerboard noise pattern. Other telltale signs of deepfakes include inconsistencies in color or tone, oversaturation or undersaturation of colors, and unnatural image noise patterns. See the top row in Figure 26. Fluorescent bleed, where bright colors bleed onto the hair or face of a person in the image from the background, is also a potential indicator of a deepfake (the bottom row in Figure 26). The human attention system is naturally adept at quickly recognizing these patterns, making them useful tools for identifying generated images.
Figure 26: Top row: problems with color and noise in generated images. Bottom row: fluorescent colors sometimes bleed in from background onto the hair or face.
Images with Cartoonish Look. AI generated images may look cartoonish or may look like a painting. This could be due to several reasons such as inconsistent or unnatural image texture, lack of depth, or focus. Some examples are shown in Figure27.
Figure 27: Some generated images that look cartoonish or look like paintings.
Technical Details. AI-generated images may contain technical details that are either incorrect or appear as random shapes. For example, furniture legs can be particularly challenging for AI to accurately render, resulting in incorrect numbers of legs or physically impossible configurations. These issues can be attributed to the inherent difficulty of modeling complex objects and the limitations of the AI’s understanding of real-world. Some example failures are shown in Figure 28.
Figure 28: Generated images with flawed details.
Accurately rendering all details in complex scenes or crowd scenes, such as those depicted in Figures 29 and 30, can be particularly challenging for AI. The complexity of these scenes makes it difficult for the AI to accurately model every detail and can lead to errors in object placement, lighting, perspective, and other features. Despite the challenges, AI technology continues to improve, and advancements are being made in the generation of more realistic and believable large and crowd scenes.
Figure 29: Example failures of generated complex scenes. Achieving accurate and detailed rendering in these types of images is particularly difficult due to the large number of objects and the intricate relationships between them.
Figure 30: Generated crowd scenes with issues.
4 Discussion
4.1 Other Cues
In addition to the cues discussed above, there are several other indicators that can be used to identify deep fakes. One such method involves examining the metadata of an image or conducting a reverse Google search to verify its authenticity. Additionally, common sense can be applied to detect images that are likely to be generated, such as a shark swimming down a street or aliens eating sushi in a Chinese restaurant. Other indications of a deep fake include lack of motion blur, unnatural bokeh, all objects appearing in focus, and repeated patterns in the image.
4.2 Some Challenging Objects
Generative models face particular challenges when it comes to generating images of objects such as clocks, Lego houses, chessboards, carpets, circuit boards, basketballs, glasses of water, dice, diagrams and tables, keyboards, and computer screens. One of the reasons for this is that these types of images contain many repeated patterns, which can be difficult for the model to accurately capture. Several examples of failed attempts to generate these objects can be seen in Figures 31 and 32. This list of challenging objects can be used to assess and compare the performance of different image generation models.
Figure 31: Some object that are difficult for models to generate.
Figure 32: Additional challenging objects for models to generate.
4.3 Memorization and Copyright
As previously mentioned, a method for identifying whether an image is generated or not is through reverse image search. Generative models may memorize images partially or in their entirety, as seen in the examples presented in Figure 33. This phenomenon has raised concerns regarding copyright infringement, as generated images may include watermarks from the original images. For more information on this issue, please refer to the this link.
Figure 33: The images on the left side of each pair are generated by StableDiffusion. One pair shows an oil painting of American Gothic by Hieronymus Bosch, while the other pair depicts The Ghosts of Hokusai.
Figure 34: Images that violate copyright generated by StableDiffusion.
5 Conclusion and Future Work
This paper lists several qualitative indicators for identifying deepfakes. These indicators not only enable us to address the issue of fake images but also underscore the differences between generated and real-world content [7]. Furthermore, they serve as a checklist for evaluating image generation models.
It should be noted that as algorithms improve, some of these clues may become obsolete over time. However, this does not mean that these models will not make any of these mistakes in generating images. It may be necessary to use a combination of these indicators to identify deepfakes, as there is no one-size-fits-all solution.
Image generation models are becoming increasingly widespread and accessible. However, in the wrong hands, these algorithms can be used to create propaganda and other forms of fake media. In a world rife with fake news [17], we have learned not to believe everything we read. Now, we must also exercise caution when it comes to visual media. The blurring of lines between reality and fiction could transform our cultural landscape from one primarily based on truth to one characterized by artificiality and deception. As we have demonstrated with the set of cues presented here, it is possible to identify fake images. In fact, we were able to use some of these indicators to detect fake faces with high accuracy in the quiz available on whichfaceisreal.com.
While visual inspection can be a valuable tool for identifying deepfakes, it may not be sufficient to detect all types of generated images. Therefore, incorporating other methods such as machine learning algorithms or forensic analysis can offer a more thorough approach to deepfake detection. Additionally, staying up-to-date of the latest techniques and technologies in this field is crucial, as it is continually advancing.
In this study, we focused on still images. However, for videos, additional indicators beyond those outlined here, such as motion and optical flow, as well as the synchronization of lip, face, and head movements over time, can also be significant factors [3]. Educating individuals on the cues outlined in this paper may aid in combatting deepfake proliferation. It would be worthwhile to investigate whether individuals can be effectively trained to become experts in this area.